In 28 years, Mark will be 5times as old as he is now. Find his present age.
Answer by Maths68(1474)

Let
Present age of Mark = m
In 28 years means we have to add 28 to his present age
m+28
then
he will be 5 times as old as he is now
5m=m+28
5m-m=28
4m=28
Divide by 4 both sides of the equation
4m/d=28/4
m=7
Present age of Mark = m = 7 years old